please help me solve this equation
3+2x sqrt3=5

Answer by mananth(16946)

3+2x sqrt3=5
2xsqrt3= 5-2
4x^2*3 =9
12x^2=9
x^2= 9/12
x^2= 3/4
x= sqrt3 /2
.